OpenStudy (anonymous):

write the equation in slope intercept form of the line that passes through (0,4) and is perpendicular to y = 2/3x-3

OpenStudy (aroub):

y=mx+b , m= -3/2 , (0,4) 4=-3/2(0)+b 4=b y=-3/2+4

hero (hero):

Hi cute aroub

OpenStudy (aroub):

lol u liked my pic?!

hero (hero):

Yes,it is very nice

OpenStudy (anonymous):

can you move this to someone else question

hero (hero):

I was just saying Hi. Calm down

hero (hero):

Looks like she answered your question already

OpenStudy (anonymous):

why say hi to someone you dont know? haha and yes i realize that and im very greatful for that

hero (hero):

I say Hi to people I don't know all the time. It's called being friendly

hero (hero):

If you don't say Hi to people you don't know, you'll never gain any new friends.

OpenStudy (aroub):

HERO, heheh thank you..!! :D / prettyboy, no problem! :)